# Latent Self-Consistency for Reliable Majority-Set Selection in Short- and Long-Answer Reasoning

### v1

@@ Abstract
We introduce Latent Self-Consistency (LSC), which selects the most semantically consistent response using learnable token embeddings. A lightweight forward generation of summary tokens increases inference time by less than 1% and requires no changes to the model architecture.
@@
Across 6 short-form and 5 long-form reasoning benchmarks (e.g., MATH, MMLU, TruthfulQA), LSC surpasses SC, USC and WUCS on all short-form and long-form ones on average, while maintaining negligible computational overhead. These results position LSC as a practical consistency-selection method that works reliably across answer formats. Additionally, LSC provides well-calibrated confidence estimates, maintaining low Expected Calibration Error across both answer formats.
1 Introduction
@@
Large Language Models (LLMs) have achieved remarkable success across diverse natural language processing tasks, from mathematical reasoning to code generation Brown et al. (2020); Chowdhery et al. (2023); OpenAI (2023). However, as inherently probabilistic generators, they frequently yield divergent outputs on repeated runs, undermining consistency and reliability in downstream applications. To mitigate this variability, Self-Consistency (SC) Wang et al. (2022) generates multiple candidate responses through independent sampling and applies majority voting to select the most frequent answer. This mechanism has proven highly effective, achieving substantial accuracy improvements in mathematical reasoning and multiple-choice benchmarks by filtering out anomalous reasoning paths.
@@
However, SC’s reliance on exact string matching severely limits its applicability. While effective for "short-answer" tasks such as numerical problems or multiple-choice questions where alternative surface forms are rare, SC fails in "long-answer" scenarios including code generation, summarization, or detailed explanations, where semantically equivalent responses often differ lexically and structurally. To address this limitation, recent consistency-based extensions have emerged. Universal Self-Consistency (USC) Chen et al. (2024) employs the LLM itself as a semantic judge by concatenating all $N$ candidates into a single prompt for evaluation; however, this approach incurs substantial computational overhead (approximately 10% additional inference time) and significant memory usage (around 15% additional memory), while introducing potential judge-model biases. Weighted Unigram Consistency Score (WUCS) Jain et al. (2024) offers a lightweight alternative using term-frequency weighted Jaccard similarity with probabilistic weighting, but sacrifices accuracy on short-answer tasks and fails to capture semantic coherence in complex reasoning chains.
@@
In this work, we introduce Latent Self-Consistency (LSC), a novel framework that achieves the best of both worlds: matching SC’s high accuracy on short-answer tasks while extending robustly to long-answer scenarios, all with minimal computational overhead (less than 1% additional inference time and negligible memory usage). LSC appends a small set of trainable tokens after the EOS token of each generated response and extracts their final-layer representations, enabling efficient KV cache reuse, to obtain compact embeddings that capture the semantic essence of each response. By computing pairwise cosine similarities in this learned embedding space, LSC identifies the most consistent response with negligible computational overhead and no complex text processing.
Our contributions are fourfold:
@@
  • Universal majority selection across answer formats. We present the first consistency-based selection method that achieves competitive or superior performance on both short-answer tasks (matching SC) and long-answer tasks (surpassing USC and WUCS), without specific mode change per answer format.
@@
  • Efficient semantic consistency via learned suffix embeddings. By encoding response semantics into compact learned embeddings, LSC achieves USC-level semantic understanding with significantly reduced latency and memory overhead, making it practical for real-time deployment.
@@
  • Reliable and calibrated confidence estimates across tasks. LSC enables trustworthy uncertainty quantification, achieving low Expected Calibration Error (ECE) scores for both short-answer and long-answer benchmarks.
@@
  • Dynamic boundary detection for robust majority set identification. We introduce a dynamic Top-$K$ boundary detection algorithm that automatically filters out noisy outliers, further enhancing selection accuracy by focusing on the most semantically coherent subset of responses.
Figure 1: Overview of the Latent Self-Consistency (LSC) inference framework. Step 1: Multiple candidate responses are generated. Step 2: A sequence of summary tokens is appended to each response, and a lightweight pass through the LLM yields the last hidden states of these tokens, which are averaged to produce a single representation per response. Step 3: Pairwise cosine similarities between these response representations are computed, and the candidate with the highest mean cosine similarity to all others is selected as the most consistent output.
2 Related Works
2.1 Consistency-Based Selection in LLMs
@@
Self-Consistency (SC) Wang et al. (2023) generates $N$ independent samples and selects the most frequent answer via exact string matching, yielding strong gains on "short answer form" tasks but failing when semantically equivalent outputs vary lexically. Universal Self-Consistency (USC) Chen et al. (2024) uses the LLM itself as a referee by concatenating all $N$ candidates into a single prompt and performing one additional forward pass. However, this approach requires constructing entirely new prompts, preventing KV cache reuse and introducing substantial latency and potential judge-model bias. Unigram Consistency Score (UCS) and its weighted variant WUCS Jain et al. (2024) compute probability-weighted overlap measures for text similarity assessment, which while efficient, struggle to capture deep chain-of-thought coherence. Fine-grained Self-Consistency (FSC) Wang et al. (2024) integrates segment-level consensus by re-generating and synthesizing responses from scratch, incurring extra latency due to these additional generative passes. In contrast, our Latent Self-Consistency (LSC) extracts small, trainable token embeddings by simply appending learnable tokens after the EOS token of each generated response, enabling KV cache reuse for highly efficient forward passes. This approach unifies SC’s accuracy on short-answer tasks with robust, low-latency selection for long-answer tasks.
2.2 Prompt Tuning and Learnable Token Embeddings
@@
Existing prompt tuning approaches Lester et al. (2021); Li and Liang (2021) typically append learnable tokens to user prompts for domain adaptation without full model fine-tuning. These methods operate in a pre-generation manner, modifying the input before the model generates responses.
@@
Our work introduces the first approach to learn semantic summary representations of generated responses through learnable token embeddings. Our LSC method operates in a post-generation manner, appending learnable tokens after the assistant’s response generation is complete to capture the semantic essence of each response. We refer to these as "summary tokens" since they are trained to encode compact semantic summaries of the full responses. Crucially, since we only update the summary token embeddings while keeping the base LLM frozen, our method preserves the model’s original natural language capabilities without any degradation. This design enables efficient semantic consistency evaluation through lightweight forward passes while maintaining complete separation between consistency measurement and the model’s core competencies.
3 Method
Overview. Figure 1 illustrates the Latent Self-Consistency (LSC) inference pipeline in three sequential phases. Step 1, candidate generation: we prompt the LLM to produce $N$ responses. Step 2, post-generation encoding: we append $K$ learnable summary tokens after EoS token to each response and perform a lightweight forward pass through the LLM to yield final-layer representations of these tokens; these representations are averaged and normalized to form one compact representation per response. Step 3, majority answer detection: pairwise cosine similarities among representations are computed to identify and select the most representative response that aligns with the semantic majority.
We now describe the implementation of this pipeline through three key components: training procedure for learning semantic representations, efficient inference procedure, and majority answer selection.
3.1 Training Phase
3.1.1 Dataset Construction
@@
We assemble a balanced mix of short-answer and long-answer samples using LLM-based pseudo-labeling. For each question, an LLM generates an answer from which we extract the answer string as the label, identifying which responses converge on the same conclusion. Short-answer data are drawn from GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021b), and TriviaQA Joshi et al. (2017). Long-answer samples are derived from MMLU Hendrycks et al. (2021a) by prompting chain-of-thought rationales and removing the final option letter to yield free-form explanations with known pseudo-labels. We maintain a 50:50 ratio of short to long formats. Detailed dataset construction procedures, statistics, and examples are provided in Appendix A, with prompt templates in Appendix B.
3.1.2 Learning representations via Supervised Contrastive Learning
@@
To bring semantically consistent responses closer together, we use supervised contrastive learning: embeddings of responses sharing the same answer label are pulled together, while those with different labels are pushed apart. Concretely, we append $K$ learnable special tokens after each response and fine-tune only their embeddings (e.g. $K=6$, $d=4096$ for 8B LLM, updating $3\times 10^{-6}$ of parameters). Let $\{h_{i,1},\dots,h_{i,K}\}$ be the final-layer representations of those tokens for response $i$; we compute
@@
With $y_{i}$ the answer label, the supervised contrastive loss He et al. (2020); Khosla et al. (2020) over a batch of size $N$ is
@@
3.2 Inference Phase
At inference, each candidate is appended with the same $K$ tokens after the EoS token and encoded in one lightweight forward pass to produce representations $\{z_{i}\}$. Unlike USC which concatenates all responses into new prompts requiring full re-computation, LSC appends tokens directly after the EoS token of each generated response, enabling reuse of the existing key-value (KV) cache from the generation phase. This means the forward pass only needs to compute representations for the $K$ additional tokens rather than processing entire sequences from scratch, resulting in minimal computational overhead (less than 1% latency increase and negligible memory usage). We then compute the cosine similarity matrix
$S_{ij}=\frac{z_{i}\cdot z_{j}}{\|z_{i}\|\|z_{j}\|}.$
3.3 Majority Answer Selection Phase
Our goal is to select the majority answer by identifying the response that has the highest average similarity to all other responses, effectively choosing the response that best represents the semantic consensus.
3.3.1 Exponentially-Weighted Mean Selection
We score each response by
$w_{i}=\frac{1}{N-1}\sum_{j\neq i}\exp\bigl{(}S_{ij}/\tau^{\prime}\bigr{)},$
with $\tau^{\prime}=0.5$. This exponentially-weighted aggregation is crucial for robust majority answer selection. A simple arithmetic mean would treat all responses equally, causing the selection to be pulled toward the centroid of all responses rather than focusing on the true majority. This centroid-pulling effect is particularly problematic when outlier responses exist. By exponentially weighting similarities, we dramatically reduce the influence of low-similarity responses (outliers), ensuring that the selection is dominated by responses that are semantically consistent with the majority. This weighting scheme effectively downweights inconsistent outliers while amplifying the influence of responses that align with the semantic consensus.
Short Long Increased Extra-resources
Models Methods Average(%) Average(%) Time(%) Memory(%)
LLaMA3.1- 8B-Instruct SC 72.2 N/A 0.0 0.0
WUCS 70.0 63.8 0.0 0.0
USC 67.7 64.8 7.4 16.2
LSC(Ours) 72.3 65.2 0.2 0.0
   +Dynamic TopK 72.8 65.8 0.2 0.0
Qwen3-8B SC 77.1 N/A 0.0 0.0
WUCS 74.9 75.4 0.0 0.0
USC 75.4 74.7 18.2 20.2
LSC(Ours) 76.8 74.9 0.6 0.0
   +Dynamic TopK 76.5 74.9 0.6 0.0
LLaMA3.3- 70B-Instruct SC 82.7 N/A 0.0 0.0
WUCS 81.1 76.7 0.0 0.0
USC 82.5 76.8 9.7 7.5
LSC(Ours) 82.6 77.6 0.9 0.0
   +Dynamic TopK 82.7 77.6 0.9 0.0
Table 1: Overall results showing LSC’s superior balance of performance and efficiency. Time/memory columns show the percentage overhead added to baseline generation time when selecting from 10 candidates.
3.3.2 Dynamic Top-$K$ for Noise Reduction
To minimize the influence of noisy outliers in majority answer selection, we implement a dynamic boundary detection algorithm that automatically determines the optimal majority subset size across all responses. This approach excludes outlier responses from interfering with accurate majority identification.
For each candidate value of $K\geq 2$, we first compute the mean similarity score for every response $i$ over its top-$K$ most similar peers:
$w_{i}^{(K)}=\frac{1}{K}\sum_{j=1}^{K}S_{i,\sigma_{i}(j)},$
where $\sigma_{i}$ sorts other responses by their similarity to response $i$ such that $S_{i,\sigma_{i}(1)}\geq S_{i,\sigma_{i}(2)}\geq\cdots$.
Next, for each $K$, we identify the response with the highest score:
$w^{(K)}_{\max}=\max_{i}w_{i}^{(K)}.$
We then compute the drop in maximum score as $K$ increases:
$\Delta^{(K)}=w^{(K-1)}_{\max}-w^{(K)}_{\max}$.
The optimal majority subset size is determined as $K^{*}=\arg\max_{K}\Delta^{(K)}-1$, selecting the value immediately before the largest drop in maximum similarity. This identifies the natural boundary where including additional responses would significantly introduce noise from inconsistent outliers. Finally, we select the response with the highest score at this optimal size: $\arg\max_{i}w_{i}^{(K^{*})}$.
Figure 2: Consistency (i.e., the fraction of counts the selected response belongs to the true majority set) versus Majority set size on the MATH dataset using the LLaMA3.1-8B-Instruct model. At low majority set size levels, LSC (ours) achieves substantially higher consistency than all other methods, and incorporating Dynamic Top-K further boosts consistency in this regime.
In-domain Out-of-domain Average Models Decoding Methods GSM8K MATH TriviaQA MMLU TruthfulQA MC1 CommonsenseQA LLaMA3.1 8B-Inst. Vanilla (1 path) 83.2 39.4 62.4 68.8 57.2 74.9 64.3 Self Consistency 92.2 52.5 72.7 73.6 63.0 79.4 72.2 WUCS 89.8 45.6 73.2 72.8 61.8 76.6 70.0 USC 86.2 44.8 70.1 71.1 59.6 74.6 67.7 LSC (Ours) 92.3 52.6 74.1 73.8 62.1 79.0 72.3 +Dynamic TopK 92.2 52.9 75.7 73.9 62.7 79.1 72.8 Qwen3-8B Vanilla (1 path) 92.8 71.9 48.1 76.7 72.6 81.6 74.0 Self Consistency 94.5 76.7 56.1 77.2 75.2 82.6 77.1 WUCS 93.8 72.1 49.9 76.8 75.0 81.9 74.9 USC 93.7 75.6 50.9 76.0 74.4 81.8 75.4 LSC (Ours) 94.5 75.9 55.5 77.1 74.9 82.7 76.8 +Dynamic TopK 94.4 76.6 53.2 77.0 75.4 82.6 76.5 LLaMA3.3 70B-Inst. Vanilla (1 path) 95.9 65.1 85.2 84.1 73.9 84.4 81.4 Self Consistency 96.9 68.6 86.3 84.6 75.0 84.7 82.7 WUCS 95.5 63.7 84.6 84.3 74.2 84.1 81.1 USC 95.8 68.2 86.5 84.6 74.8 84.9 82.5 LSC (Ours) 96.9 68.7 86.3 84.2 75.3 84.3 82.6 +Dynamic TopK 97.0 68.7 86.4 84.2 75.3 84.5 82.7
Table 2: All listed datasets produce short-answer outputs. In-domain sets were used for summary-token training; Out-of-domain sets were not. The highest score in each column is in bold, and the second-highest is underlined. All scores are reported as percentages.
4 Experiments
4.1 Experimental Setup
Short-Answer Tasks: We use 6 benchmarks including mathematical reasoning (GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021b)), factual knowledge (TriviaQA Joshi et al. (2017), MMLU Hendrycks et al. (2021a)), and commonsense reasoning (CommonsenseQA Talmor et al. (2018), TruthfulQA-MC1 Lin et al. (2022)). The first four datasets are considered in-domain (used during representations learning), while the latter two are out-of-domain.
Long-Answer Tasks: We evaluate on 5 benchmarks requiring extended responses: open-ended QA (TruthfulQA Lin et al. (2022), MSMARCO-NLG Bajaj et al. (2016)), code generation (HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and their enhanced versions HumanEval+, MBPP+ Liu et al. (2023)), and summarization (CNN/DailyMail Hermann et al. (2015)).
Setup: For each dataset, we sample up to 1,000 examples and generate 10 candidate responses per question using nucleus sampling (top-$p=0.95$, temperature 0.9). We evaluate three instruction-tuned models: LLaMA3.1-8B-Instruct, LLaMA3.3-70B-Instruct Dubey et al. (2024), and Qwen3-8B Bai et al. (2023). All models use six learnable summary tokens (<|Summary1|>, …, <|Summary6|>) trained on a balanced mixture of short and long-answer samples.
4.2 Results
Models Decoding Methods MSMARCO v2.1 Natural Language Generation TruthfulQA Semantic Similarity BLEU 1 BLEU 2 BLEU 3 BLEU 4 ROUGE L Truth(%) LLaMA3.1- 8B-Instruct Vanilla (1 path) 0.8805 0.4122 0.3137 0.2488 0.2005 0.5075 39.9 WUCS 0.8815 0.3904 0.2996 0.2392 0.1938 0.4987 42.8 USC 0.8812 0.4065 0.3103 0.2463 0.1989 0.5086 44.2 LSC(Ours) 0.8838 0.4269 0.3274 0.2613 0.2118 0.5207 43.5    +Dynamic TopK 0.8861 0.4336 0.3333 0.2667 0.2167 0.5270 44.4 Qwen3-8B Vanilla (1 path) 0.8842 0.3750 0.2876 0.2303 0.1875 0.4947 43.6 WUCS 0.8835 0.3620 0.2788 0.2242 0.1835 0.4886 44.7 USC 0.8824 0.3608 0.2772 0.2223 0.1812 0.4871 43.1 LSC(Ours) 0.8838 0.3756 0.2896 0.2330 0.1906 0.4964 44.9    +Dynamic TopK 0.8836 0.3761 0.2897 0.2330 0.1905 0.4958 44.6 LLaMA3.3- 70B-Instruct Vanilla (1 path) 0.8926 0.4694 0.3662 0.2981 0.2463 0.5618 48.0 WUCS 0.8914 0.4277 0.3343 0.2716 0.2242 0.5500 48.7 USC 0.8913 0.4623 0.3603 0.2930 0.2423 0.5596 50.6 LSC(Ours) 0.8918 0.4724 0.3698 0.3016 0.2499 0.5629 48.8    +Dynamic TopK 0.8922 0.4768 0.3733 0.3042 0.2519 0.5652 49.3

### v2

# Latent Self-Consistency for Reliable Majority‑Set Selection in Short‑ and Long‑Answer Reasoning

@@ Abstract
We introduce Latent Self-Consistency (LSC), which selects the most semantically consistent response using learnable token embeddings. LSC’s lightweight forward processing of summary tokens only introduces negligible runtime overhead (at most $0.9\%$) on top of standard decoding of the base LLM, and requires no changes to the model architecture.
@@
Across 6 short-form and 5 long-form reasoning benchmarks (e.g., MATH, MMLU, TruthfulQA), LSC surpasses SC, USC, and WUCS on both short-form and long-form on average performance, while adding negligible computational overhead on vanilla inference. These results position LSC as a reliable consistency-selection method that works effectively across various answer formats. Additionally, LSC provides well-calibrated confidence estimates, maintaining low expected calibration error across both answer formats.
Codehttps://github.com/jeongseokO/LatentSC˙official
Datasetshttps://huggingface.co/jeongseokoh
Introduction
@@
Large Language Models (LLMs) have achieved remarkable success across diverse natural language processing tasks, from mathematical reasoning to code generation (Brown et al. 2020; Chowdhery et al. 2023; OpenAI 2023). However, as inherently probabilistic generators, they frequently yield divergent outputs on repeated runs, undermining consistency and reliability in downstream applications. To mitigate this variability, Self-Consistency (SC) (Wang et al. 2023) generates multiple candidate responses through independent sampling and applies majority voting to select the most frequent answer. This mechanism has proven highly effective, achieving substantial accuracy improvements in mathematical reasoning and multiple-choice benchmarks by filtering out anomalous reasoning paths.
@@
SC’s reliance on exact string matching, however, severely limits its applicability. While effective for “short-answer” tasks such as numerical problems or multiple-choice questions where alternative surface forms are rare, SC fails in “long-answer” scenarios including code generation, summarization, or detailed explanations, where semantically equivalent responses often differ lexically and structurally. To address this limitation, several consistency-based extensions have emerged. Universal Self-Consistency (USC) (Chen et al. 2024) employs the LLM itself as a semantic judge by concatenating all candidates into a single prompt for evaluation; however, this approach incurs substantial computational overhead (approximately 10% additional inference time) and significant memory usage (around 15% additional memory), while introducing potential judge-model biases. Weighted Unigram Consistency Score (WUCS) (Jain et al. 2024) offers a lightweight alternative using term-frequency–weighted Jaccard similarity with probabilistic weighting, but sacrifices accuracy on short-answer tasks and fails to capture semantic coherence in complex reasoning. These alternative methods sometimes achieve competitive accuracy, but such gains only partly arise from putting together truly consistent reasoning paths. For example, USC often relies on LLM reasoning when it selects a more accurate response that disagrees with the majority response.
In practice, this leads to an unsatisfying trade-off: SC performs strongly on short-answer benchmarks but breaks down on long-form tasks, whereas USC and WUCS are better suited to free-form generation yet often degrade performance on short-answer evaluation. Empirically, our experiments (Appendix K) show that, given only the input question, it is difficult to reliably predict whether the model will produce a short-answer–style or a long-answer–style response, making it impractical to route queries to different decoding schemes based on an assumed answer format.
Figure 1: Overview of the Latent Self-Consistency (LSC) inference framework.
@@
In this work, we introduce Latent Self-Consistency (LSC), a framework that aims to achieve the best of both worlds: matching SC’s high accuracy on short-answer tasks while extending robustly to long-answer scenarios, with minimal additional computational cost. LSC appends a small set of trainable summary tokens after the EOS token of each generated response and reuses the key–value (KV) cache from the original generation to compute their final-layer representations, yielding compact embeddings that capture the semantic essence of each response. LSC introduces a negligible runtime overhead (at most 0.9%) compared to standard decoding and requires no changes to the model architecture. By computing pairwise cosine similarities with these learned embeddings, LSC identifies the most consistent response without constructing new prompts or performing any complex text processing.
Our contributions are as follows:
@@
  • Universal majority selection across answer formats. We present the first consistency-based selection method that delivers strong performance on both short- and long-answer tasks while using a single unified approach, with no format-dependent tuning or heuristics.
@@
  • Efficient semantic consistency via learned suffix embeddings. By encoding response semantics into compact learned embeddings obtained from a few summary tokens, LSC attains USC-level semantic discrimination with substantially reduced runtime and memory overhead, making it practical for real-time deployment.
@@
  • Reliable and calibrated confidence estimates across tasks. LSC enables uncertainty quantification, achieving competitive expected calibration error (ECE) scores for both short- and long-answer benchmarks.
@@
  • Dynamic boundary detection for robust majority set identification. We introduce a dynamic Top-$K$ boundary detection algorithm that automatically filters out noisy outliers, further enhancing selection accuracy by focusing on the most semantically coherent subset.
Related Works
Consistency‐Based Selection in LLMs
@@
Self‐Consistency (SC) (Wang et al. 2023) generates $N$ independent samples and selects the most frequent answer via exact string matching, yielding strong gains on ”short answer form” tasks but failing when semantically equivalent outputs vary lexically. Universal Self‐Consistency (USC) (Chen et al. 2024) uses the LLM itself as a referee by concatenating all $N$ candidates into a single prompt and performing one additional forward pass. However, this approach requires constructing entirely new prompts, preventing KV cache reuse and introducing substantial latency and potential judge-model bias. Unigram Consistency Score (UCS) and its weighted variant WUCS (Jain et al. 2024) compute probability-weighted overlap measures for text similarity assessment, which while efficient, struggle to capture deep chain-of-thought coherence. Fine‐grained Self‐Consistency (FSC) (Wang et al. 2024) integrates segment-level consensus by re-generating and synthesizing responses from scratch, incurring extra latency due to these additional generative passes. In contrast, our LSC extracts small, trainable token embeddings by simply appending learnable tokens after the EOS token of each generated response, enabling KV cache reuse for efficient forward passes. This approach unifies SC’s accuracy on short-answer tasks with robust, low-latency selection for long-answer tasks.
Short Long Overhead on top of Vanilla
Models Methods Average(%) Average(%) Time(%) Memory(%)
LLaMA3.1‑ 8B-Instruct SC 72.2 N/A 2E-3 0.0
WUCS 70.0 63.8 3E-2 0.0
USC 67.7 64.8 7.4 16.2
LSC(Ours) 72.3 65.2 0.2 5E-3
   +Dynamic TopK 72.8 65.8 0.2 5E-3
Qwen3-8B SC 77.1 N/A 1E-3 0.0
WUCS 74.9 75.4 2E-2 0.0
USC 75.4 74.7 18.2 20.2
LSC(Ours) 76.8 74.9 0.6 5E-3
   +Dynamic TopK 76.5 74.9 0.6 5E-3
LLaMA3.3‑ 70B-Instruct SC 82.7 N/A 1E-4 0.0
WUCS 81.1 76.7 8E-3 0.0
USC 82.5 76.8 9.7 7.5
LSC(Ours) 82.6 77.6 0.9 1E-3
   +Dynamic TopK 82.7 77.6 0.9 1E-3
Table 1: Overall performance–efficiency trade-off of LSC. Time/Memory columns report percentage overhead over vanilla inference when generating 10 candidates.
Prompt Tuning and Learnable Token Embeddings
@@
Existing prompt tuning approaches (Lester, Al-Rfou, and Constant 2021; Li and Liang 2021) typically append learnable tokens to user prompts for domain adaptation without full model fine-tuning. These methods operate in a pre-generation manner, modifying the input before the model generates responses.
@@
Our work introduces the first approach to learn semantic summary representations of generated responses through learnable token embeddings. Our LSC method operates in a post-generation manner, appending learnable tokens after the assistant’s response generation is complete to capture the semantic essence of each response. We refer to these as ”summary tokens” since they are trained to encode compact semantic summaries of the full responses. Crucially, since we only update the summary token embeddings while keeping the base LLM frozen, our method preserves the model’s original natural language capabilities without any degradation. This design enables efficient semantic consistency evaluation through lightweight forward passes while maintaining complete separation between consistency measurement and the model’s core competencies.
Method
In-domain Out-of-domain Average
Models Decoding Methods GSM8K MATH TriviaQA MMLU TruthfulQA MC1 CommonsenseQA
LLaMA3.1 8B-Inst. Vanilla (1 path) 83.2 39.4 62.4 68.8 57.2 74.9 64.3
Self Consistency 92.2 52.5 72.7 73.6 63.0 79.4 72.2
WUCS 89.8 45.6 73.2 72.8 61.8 76.6 70.0
USC 86.2 44.8 70.1 71.1 59.6 74.6 67.7
LSC (Ours) 92.3 52.6 74.1 73.8 62.1 79.0 72.3
+Dynamic TopK 92.2 52.9 75.7 73.9 62.7 79.1 72.8
Qwen3-8B Vanilla (1 path) 92.8 71.9 48.1 76.7 72.6 81.6 74.0
Self Consistency 94.5 76.7 56.1 77.2 75.2 82.6 77.1
WUCS 93.8 72.1 49.9 76.8 75.0 81.9 74.9
USC 93.7 75.6 50.9 76.0 74.4 81.8 75.4
LSC (Ours) 94.5 75.9 55.5 77.1 74.9 82.7 76.8
+Dynamic TopK 94.4 76.6 53.2 77.0 75.4 82.6 76.5
LLaMA3.3 70B-Inst. Vanilla (1 path) 95.9 65.1 85.2 84.1 73.9 84.4 81.4
Self Consistency 96.9 68.6 86.3 84.6 75.0 84.7 82.7
WUCS 95.5 63.7 84.6 84.3 74.2 84.1 81.1
USC 95.8 68.2 86.5 84.6 74.8 84.9 82.5
LSC (Ours) 96.9 68.7 86.3 84.2 75.3 84.3 82.6
+Dynamic TopK 97.0 68.7 86.4 84.2 75.3 84.5 82.7
Table 2: Performance comparisons on short-answer benchmarks. Only In-domain sets were used for summary-token training. The highest score in each column is in bold, and the second-highest is underlined. All scores are reported as percentages.
Inference Phase
Figure 1 illustrates the Latent Self-Consistency (LSC) inference pipeline in three phases: Candidate generation, Post-generation encoding, and Majority answer detection.
For (1) Candidate generation, given an input question we prompt the base LLM and sample $N$ candidate responses. For (2) Post-generation encoding, we append $K$ learnable summary tokens at the end of each response and run a lightweight forward pass to produce response representations $\{z_{i}\}$. Unlike USC, which concatenates all responses into a new prompt and recomputes the entire sequence, LSC appends only a few tokens at the end of each candidate and reuses the existing key–value (KV) cache from the candidate generation phase. As a result, the model needs to compute representations only for the $K$ additional tokens rather than processing entire sequences from scratch. This makes LSC highly efficient as it only adds negligible overhead on top of base LM: at most $0.9\%$ runtime and 0.0050.005% memory overhead in our experiments.
For (3) Majority answer detection, we first compute the cosine similarity matrix between all pairs of response embeddings, $S_{ij}=\frac{z_{i}\cdot z_{j}}{\|z_{i}\|\|z_{j}\|}$. We then assign an exponentially weighted similarity score to each response $i$,
$w_{i}=\frac{1}{N-1}\sum_{j\neq i}\exp\bigl(S_{ij}/\tau^{\prime}\bigr),$
with $\tau^{\prime}=0.5$. A simple arithmetic mean of $\{S_{ij}\}_{j\neq i}$ treats all the responses equally and will likely pull $w_{i}$ toward the global centroid, especially in the presence of outliers. In contrast, the exponential weighting effectively suppresses low-similarity responses, resulting in $w_{i}$ being primarily determined by candidates from the majority set.
To further suppress the influence of residual outliers, we additionally propose a Dynamic Top-$K$ scheme that estimates the size of the majority set directly from the similarity structure. For each $K\geq 2$, let $\sigma_{i}$ sort other responses in decreasing similarity to $i$, and define the average similarity to the $K$ nearest neighbors
$w_{i}^{(K)}=\frac{1}{K}\sum_{j=1}^{K}S_{i,\sigma_{i}(j)},$
with the best score at that neighborhood size
$w^{(K)}_{\max}=\max_{i}w_{i}^{(K)}.$
We then track the drop between $K$ and K1K-1,
$\Delta^{(K)}=w^{(K-1)}_{\max}-w^{(K)}_{\max}$,
and choose
$K^{*}=\arg\max_{K}\Delta^{(K)}-1,$
which corresponds to the neighborhood size just before the largest decrease in maximum similarity. Finally, we select the index
$i^{*}=\arg\max_{i}w_{i}^{(K^{*})}$
as the majority choice.
Training Phase
Dataset Construction
@@
We assemble a balanced mix of short‐answer and long‐answer samples using LLM‐based pseudo‐labeling. For each question, an LLM generates an answer from which we extract the answer string as the label, identifying which responses converge on the same conclusion. Short‐answer data are drawn from GSM8K (Cobbe et al. 2021), MATH (Hendrycks et al. 2021b), and TriviaQA (Joshi et al. 2017). Long‐answer samples are derived from MMLU (Hendrycks et al. 2021a) by prompting chain‐of‐thought rationales and removing the short-form answer to yield ’free‐form explanations with known pseudo-labels’. We maintain a 50:50 ratio of short to long formats. Detailed dataset construction procedures, statistics, and examples are provided in Appendix A, with prompt templates in Appendix B.
Learning Representations via Supervised Contrastive Learning
@@
To bring semantically consistent responses closer together, we use supervised contrastive learning: embeddings of responses sharing the same answer label are pulled together, while those with different labels are pushed apart. Concretely, we append $K$ learnable special tokens after each response and fine‐tune only their embeddings (e.g. $K=6$, $d=4096$ for 8B LLM, updating $3\times 10^{-6}$ of parameters). Let $\{h_{i,1},\dots,h_{i,K}\}$ be the final‐layer representations of those tokens for response $i$; we compute
@@
With $y_{i}$ the answer label, the supervised contrastive loss (He et al. 2020; Khosla et al. 2020) over a batch of size $N$ is
@@
Figure 2: Consistency (fraction of examples where the selected response is in the true majority set) vs. majority-set size on MATH with LLaMA3.1-8B-Instruct, showing LSC outperforming all baselines and Dynamic Top-$K$ giving additional gains for small majority sets.
Models Decoding Methods MSMARCO v2.1 Natural Language Generation TruthfulQA
Semantic Similarity BLEU 1 BLEU 2 BLEU 3 BLEU 4 ROUGE L Truth(%)
LLaMA3.1‑ 8B-Instruct Vanilla (1 path) 0.8805 0.4122 0.3137 0.2488 0.2005 0.5075 39.9
WUCS 0.8815 0.3904 0.2996 0.2392 0.1938 0.4987 42.8
USC 0.8812 0.4065 0.3103 0.2463 0.1989 0.5086 44.2
LSC(Ours) 0.8838 0.4269 0.3274 0.2613 0.2118 0.5207 43.5
   +Dynamic TopK 0.8861 0.4336 0.3333 0.2667 0.2167 0.5270 44.4
Qwen3-8B Vanilla (1 path) 0.8842 0.3750 0.2876 0.2303 0.1875 0.4947 43.6
WUCS 0.8835 0.3620 0.2788 0.2242 0.1835 0.4886 44.7
USC 0.8824 0.3608 0.2772 0.2223 0.1812 0.4871 43.1
LSC(Ours) 0.8838 0.3756 0.2896 0.2330 0.1906 0.4964 44.9
   +Dynamic TopK 0.8836 0.3761 0.2897 0.2330 0.1905 0.4958 44.6
LLaMA3.3‑ 70B-Instruct Vanilla (1 path) 0.8926 0.4694 0.3662 0.2981 0.2463 0.5618 48.0
WUCS 0.8914 0.4277 0.3343 0.2716 0.2242 0.5500 48.7
USC 0.8913 0.4623 0.3603 0.2930 0.2423 0.5596 50.6
LSC(Ours) 0.8918 0.4724 0.3698 0.3016 0.2499 0.5629 48.8
   +Dynamic TopK 0.8922 0.4768 0.3733 0.3042 0.2519 0.5652 49.3